You have gone to a casino and placed a $5 bet in craps where the probability of winning is 2441495 If you win, the casino will give you $10. Find your expected value. O A. $0.07 O B. -$0.07 O D. $5 he Click to select your answer javascript doExercise(11);
The probability of flu symptoms for a person not receiving any treatment is 0.039. In a clinical trial o a common drug used to lower cholesterol, 35 of 803 people treated experienced flu symptoms. Assuming the drug has no effect on the likelihood of flu symptoms, estimate the probability that at least 35 people experience flu symptoms. What do these results suggest about flu symptoms as an adverse reaction to the drug? (a) P(X235) 0.123 (Round to four decimal places as needed.) (b) What does the result from part (a) suggest? O A. The drug increases the likelihood of flu symptoms because x 2 35 is highly unlikely O B. The drug has no effect on flu symptoms because x 2 35 is not highly unlikely. rhC. The drug has no effect on flu symptoms because x2 35 is highly unlikety O D. The drug increases the likelihood of flu symptoms because x 2 35 is not highly unlikely. Click to select your answer(s). 4W 5:44
Answer #1

1)

Payoff = 10-5 with p = 244/495

and      -5   with p = 1- 244/495
Expected value = 10 * 244/495   -5
= -0.07070


option B) is correct